# Byung-Kwan Lee KAIST

### v3

@@ Abstract
The remarkable success of Large Language Models (LLMs) and instruction tuning drives the evolution of Vision Language Models (VLMs) towards a versatile general-purpose model. Yet, it remains unexplored whether current VLMs genuinely possess quality object-level image understanding capabilities determined from ‘what objects are in the image?’ or ‘which object corresponds to a specified bounding box?’. Our findings reveal that the image understanding capabilities of current VLMs are strongly correlated with their zero-shot performance on vision language (VL) tasks. This suggests that prioritizing basic image understanding is crucial for VLMs to excel at VL tasks. To enhance object-level image understanding, we propose Crayon Large Language and Vision mOdel ([Uncaptioned image] CoLLaVO), which incorporates instruction tuning with Crayon Prompt as a new visual prompt tuning scheme based on panoptic color maps. Furthermore, we present a learning strategy of Dual QLoRA to preserve object-level image understanding without forgetting it during visual instruction tuning, thereby achieving a significant leap in numerous VL benchmarks in a zero-shot setting.
@@ 6 Limitations
Crayon Prompts, relying on a panoptic color map, which is an external source beyond VLMs, may be constrained by the performance of the segmentation model and its encompassing number of object classes. Despite this, we have achieved commendable scores across all zero-shot tasks. It is expected for [Uncaptioned image] CoLLaVO to further improve once it incorporates a plethora of visual prompts obtained from diverse sources like robust object classification or image captioning models (Lee et al., 2020, 2022, 2023; Kim et al., 2023c), object-centric causally human-interpretable information (Kim et al., 2021, 2023b), open object detection (Zhang et al., 2023a), visual grounding (Liu et al., 2023d; Ren et al., 2024), interactive or unsupervised segmentation (Kirillov et al., 2023; Kim et al., 2023a), and optical characteristic recognition model (Bautista and Atienza, 2022). Beyond its limitation, we believe our promising direction for crayon prompt-like visual cues surely makes VLMs further improve on image understanding for human-like AGI.

### v4

@@ Abstract
The remarkable success of Large Language Models (LLMs) and instruction tuning drives the evolution of Vision Language Models (VLMs) towards a versatile general-purpose model. Yet, it remains unexplored whether current VLMs genuinely possess quality object-level image understanding capabilities determined from ‘what objects are in the image?’ or ‘which object corresponds to a specified bounding box?’. Our findings reveal that the image understanding capabilities of current VLMs are strongly correlated with their zero-shot performance on vision language (VL) tasks. This suggests that prioritizing basic image understanding is crucial for VLMs to excel at VL tasks. To enhance object-level image understanding, we propose Crayon Large Language and Vision mOdel ([Uncaptioned image] CoLLaVO), which incorporates instruction tuning with Crayon Prompt as a new visual prompt tuning scheme based on panoptic color maps. Furthermore, we present a learning strategy of Dual QLoRA to preserve object-level image understanding without forgetting it during visual instruction tuning, thereby achieving a significant leap in numerous VL benchmarks in a zero-shot setting. Code is available in https://github.com/ByungKwanLee/CoLLaVO.
@@ 5 Discussion and Conclusion
Acknowledgments
This work was partially supported by two funds: IITP grant funded by the Korea government (MSIT) (No.2022-0-00984) and Center for Applied Research in Artificial Intelligence (CARAI) grant funded by DAPA and ADD (UD230017TD).
@@ 6 Limitations
Crayon Prompts, relying on a panoptic color map, which is an external source beyond VLMs, may be constrained by the performance of the segmentation model and its encompassing number of object classes. Despite this, we have achieved commendable scores across all zero-shot tasks. It is expected for [Uncaptioned image] CoLLaVO to further improve once it incorporates a plethora of visual prompts obtained from diverse sources like robust object classification or image captioning models (Lee et al., 2020, 2022, 2023; Kim et al., 2023c), object-centric causally human-interpretable information (Kim et al., 2021, 2023b), open object detection (Zhang et al., 2023a), visual grounding (Liu et al., 2023d; Ren et al., 2024), interactive or unsupervised segmentation (Kirillov et al., 2023; Kim et al., 2023a), optical characteristic recognition model (Bautista and Atienza, 2022), and other fascinating approaches Lee et al. (2024a, b); Park et al. (2024b, a); Kim et al. (2024). Beyond its limitation, we believe our promising direction for crayon prompt-like visual cues surely further improve on image understanding for human-like AGI.